OpenStudy (anonymous):

evaluate cot(arccos5/13)

OpenStudy (amistre64):

sin^2 + cos^2 = 1 1 + cot^2 = csc^2 cot^2 = csc^2 - 1

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

or simply draw out the triangle and solve for cot = x/y

OpenStudy (anonymous):

what does the arccos mean?

OpenStudy (amistre64):



OpenStudy (amistre64):

arccos relates a ratio to an angle ... its the reverse, or invers, of cosine

OpenStudy (amistre64):

for example: cos(45) = 1/sqrt(2) therefore 45 = arccos(1/sqrt(2))

OpenStudy (anonymous):

so arccos equals the sin

OpenStudy (amistre64):

trig (angle) = ratio angle = arctrig (ratio) if you know the ratio, you can rebuild the triangle no, sin and arccos are not equivalent function

OpenStudy (anonymous):

but would it be cot(12/13)

OpenStudy (amistre64):

you are confusing angles and ratios at the moment

OpenStudy (amistre64):